OpenStudy (anonymous):

Determine the equation of combined variation. Then solve for the missing value. x varies directly with y and inversely with z.x = 20 when y = 8 and z = 4.Find x when y = 4 and z = 8.@awkwardpanda

OpenStudy (awkwardpanda):

We have x = k*y/z for some k. Since x = 20 when y = 8 and z = 4, we know 20 = k*8/4 = 2*k, so k = 10. Therefore, x = 10*y/z. So when y = 4 and z = 8 we get x = 10*4/8 = 40/8 = 5. -anyone correct me if I'm wrong-

OpenStudy (anonymous):

We have the relations: \[\eqalign{ &x\propto y \\ &x\propto\frac{1}{z} }\] So then to put all in one equation we have: \[x=\frac{ky}{z}\] Where k is an real number and a constant. We can solve for \(k\) by plugging in \((x=20),(y=8),(z=4)\) \[\eqalign{ &x=\frac{ky}{z} \\ &20=\frac{8k}{4} \\ &20=2k \\ &k=10 \\ }\] So then the complete formula is: \[x=\frac{10y}{z}\] They give you y and z and so you can solve for x!

OpenStudy (awkwardpanda):

Wait. Nevermind, Keith got it.

OpenStudy (anonymous):

you are correct i am gann post another one @awkwardpanda

OpenStudy (awkwardpanda):

Alright..

OpenStudy (anonymous):

ok are you ready for another one @awkwardpanda